OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 40 miles per hour (mph), and the other car is traveling 20 mph. How long will it take before the cars are 300 miles apart? 2 hours 3 hours 4 hours 5 hours

OpenStudy (anonymous):

Nett speed = 40+20 =60mph 300 /60 = 5 hours :)